### v1

# The Surprising Soupability of Documents in State Space Models

Abstract
We investigate whether hidden states from Structured State Space Models (SSMs) can be merged post-hoc to support downstream reasoning. Inspired by model souping, we propose a strategy where documents are encoded independently and their representations are pooled—via simple operations like averaging—into a single context state. This approach, which we call document souping, enables modular encoding and reuse without reprocessing the full input for each query. We finetune Mamba2 models to produce soupable representations and find that they support multi-hop QA, sparse retrieval, and long-document reasoning with strong accuracy. On HotpotQA, souping ten independently encoded documents nearly matches the performance of a cross-encoder trained on the same inputs.
1 Introduction
Many real-world NLP tasks—such as multi-document question answering, scientific summarization, and legal analysis—require reasoning over entire corpora, not just individual long documents. These tasks demand flexible integration of information spread across sources, as well as the ability to update, prune, or recombine subsets of the input. Yet today’s language models remain poorly suited for this kind of modular document reasoning. Transformer-based models [21] suffer from $\mathcal{O}(L^{2})$ attention costs and memory scaling, making full-corpus encoding expensive and inflexible.
Structured State Space Models (SSMs) offer a compelling alternative. Recent architectures such as Mamba [8] and Mamba2 [6] process long sequences in linear time, compressing them into fixed-size hidden states that support efficient downstream decoding. This makes them well-suited for long-context settings. However, most current pipelines treat corpora as monolithic inputs—concatenating documents into a single sequence before encoding. This approach is brittle and inefficient: modifying even a single document requires re-encoding the entire corpus, and joint encoding eliminates opportunities for reuse across tasks or queries.
Figure 1: Computation Graphs for Corpus Encoding. Top: In traditional concatenation-based encoding, all documents $\{d_{1},\dots,d_{k}\}$, the query $q$, and answer $a$ are flattened into a single input sequence and processed end-to-end by an SSM. This requires joint re-encoding for every change to the input. Bottom: In our proposed state souping approach, each document $d_{i}$ is encoded independently by a shared SSM, producing per-document hidden states $\{h_{1},\dots,h_{k}\}$ which are pooled into a single representation $h_{\text{soup}}$ (e.g., via sum or average). This pooled state is then used, alongside the query $q$, to drive downstream prediction. The design supports parallel encoding, modular reuse, and post-hoc corpus composition.
In this work, we explore whether it is possible to modularize document representation in SSMs through a strategy inspired by model souping—a technique that averages parameters across finetuned checkpoints to combine capabilities [25]. We ask: if SSMs encode individual documents into fixed-length hidden states, can those representations be merged post-hoc—for example, by averaging—while still supporting downstream tasks like question answering? This capability would be especially valuable in retrieval-augmented generation, where cross-encoder approaches must re-encode the top-$k$ retrieved documents for every new query. In contrast, if document states are soupable, one could pre-encode the full corpus once, cache each document’s hidden state, and dynamically pool a subset at inference time—enabling scalable, query-specific reconfiguration without re-encoding.
We operationalize this idea as corpus encoding via state souping. Each document is independently encoded by a shared SSM to produce layerwise hidden states. These are then pooled using simple commutative operators (e.g., average, sum, max) into a single fixed-size representation, which conditions the decoder alongside the query. We define soupability as the property of document encodings that allows them to be pooled in this way while preserving the information needed for downstream reasoning.
Surprisingly, we find that Mamba2 encoders produce representations with strong soupability: even without joint encoding, a finetuned decoder can learn to interpret pooled document states effectively. In this setting, the encoder remains frozen, and only the decoder adapts—indicating that the individual document encodings already retain task-relevant information that can be meaningfully composed through pooling. With full encoder-decoder finetuning, performance improves further, as the encoder learns to produce even more mergeable representations.
This architecture enables a new approach to multi-document reasoning. For example, on HotpotQA [28]—a benchmark requiring multi-hop reasoning across documents—souping ten independently encoded documents achieves nearly the same accuracy as a standard cross-encoder, where all documents are jointly re-encoded with the query. More broadly, we find that corpus encoding via state souping supports single- and multi-hop QA, sparse retrieval, and long-document understanding—often matching or surpassing traditional joint encoding approaches.
Crucially, this design unlocks a new inference workflow: corpora can be encoded once, cached as hidden states, and later reassembled dynamically for downstream tasks. This supports post-hoc, retrieval-based reasoning at scale, without the need for repeated full-sequence processing.
Contributions.
We present a systematic investigation into the soupability of document representations in Structured State Space Models (SSMs) using Mamba2-2.7B [6] and Mamba2-8B [22]. Our contributions are: (1) We introduce document souping, a mechanism for merging per-document hidden states using lightweight pooling operators and injecting the result into an SSM decoder. (2) We demonstrate that, after training, Mamba2 models with soup-based representations can support multi-hop reasoning on HotpotQA, achieving performance comparable to or exceeding joint encoding. (3) We analyze soupability across diverse tasks and conditions—including long-document QA (RACE[14]/QuALITY[16]), sparse retrieval (RULER[11]/NIAH), and document count generalization—showing that souped states preserve both local and relational information. (4) We compare pooling strategies and show that soupability is robust to operator choice, with no-norm averaging working well across settings.
2 Methods
We investigate when document representations from Structured State Space Models (SSMs) can be merged post-hoc—using simple commutative operations like averaging or summation—while still preserving the information needed for downstream tasks. We refer to this property as soupability.
2.1 Background and Notation
Structured State Space Models (SSMs) are a class of sequence models that compress long inputs into fixed-size hidden representations through linear recurrence mechanisms. Unlike attention-based models, SSMs offer subquadratic compute and memory scaling, making them well suited for long-context settings.
We denote the encoder component of an SSM as a function $\mathrm{SSM}_{\theta}$, which maps a document $d$ to a sequence of layer-wise hidden states $\{h^{(1)},\dots,h^{(L)}\}$. Each $h^{(l)}\in\mathbb{R}^{d}$ summarizes the input up to layer $l$, and $L$ denotes the total number of layers. In standard pipelines, this encoder is applied to the entire input sequence jointly, and the resulting states are passed to a decoder for generation or prediction.
In this work, we leverage the layerwise structure of SSMs to explore whether documents can be encoded independently and later recombined in a modular fashion. Our central question is whether these hidden states—computed separately per document—can be pooled into a single representation suitable for downstream reasoning. We refer to this approach as corpus encoding via state souping.
2.2 Corpus Encoding via State Souping
Our proposed corpus encoding strategy is illustrated in Figure 1 and formalized in Appendix A.3. Given a set of documents $\{d_{1},\dots,d_{k}\}$ and a query $q$, each document is passed independently through a shared SSM encoder, yielding a set of hidden states $\{h_{1}^{(l)},\dots,h_{k}^{(l)}\}$ at each layer $l$. These are then combined using a commutative pooling operator—typically elementwise average, sum, or max—resulting in a single pooled state $h_{\text{soup}}^{(l)}$ for each layer.
We optionally explore unit normalization, applied either before pooling (to each $h_{i}^{(l)}$) or after pooling (to $h_{\text{soup}}^{(l)}$):
$\tilde{h}_{i}^{(l)}=\frac{h_{i}^{(l)}}{\|h_{i}^{(l)}\|},\quad h_{\mathrm{soup}}^{(l)}=\mathrm{normalize}\left(\sum_{i=1}^{k}\tilde{h}_{i}^{(l)}\right).$
Once pooled, these hidden states $\{h_{\mathrm{soup}}^{(1)},\dots,h_{\mathrm{soup}}^{(L)}\}$ are injected into the decoder alongside the query $q$. The decoder produces the answer $\hat{y}$ conditioned on the souped representation:
$\hat{y}=\mathrm{SSM}_{\theta}\left(q\;\middle|\;\{h_{\mathrm{soup}}^{(l)}\}_{l=1}^{L}\right).$
This design enables efficient parallel encoding, modular document reuse, and flexible corpus reconfiguration at inference time. Because the encoder processes each document in isolation, representations can be cached and reused across multiple queries, dramatically reducing redundant computation.
To support training of this architecture, gradients must propagate through multiple independent document encoders. Without memory optimizations, this would require storing intermediate activations for all $k$ documents, leading to memory growth linear in $k$. To address this, we apply activation checkpointing at the document level: forward activations are recomputed during the backward pass, enabling constant memory usage regardless of corpus size. This allows us to scale finetuning to wide or deep document sets efficiently.
2.3 Evaluation Dimensions
To characterize when corpus encoding via state souping is effective, we organize our analysis around two dimensions: model capacity and corpus structure.
Model capacity concerns the architectural and training properties that affect soupability. We ask: Are pretrained SSMs inherently soupable, or must they be finetuned to interpret pooled states? Does soupability improve with model size or hidden state dimensionality? And how well do models generalize across soup sizes—for example, when asked to merge more documents at test time than during training?
Corpus structure examines how input organization affects pooling success. We study whether long, contiguous documents can be segmented and recomposed via souping, or whether this method is better suited to independently authored texts. We also test whether souped representations preserve the dependencies needed for multi-hop reasoning, where answering a query requires synthesizing information from multiple documents.
Together, these questions guide our exploration of soupability as a flexible and scalable alternative to serial encoding for corpus-level reasoning in state space models.
Method Test on $2$ gold + $(n{-}2)$ distractors
2 5 10
Pretrained 8B (No Finetuned)
Concat 15.4 / 26.3 8.5 / 20.2 5.0 / 15.7
Soup w/ Average 8.7 / 12.7 2.6 / 5.0 1.7 / 3.6
Decoder-Only Finetuned 8B
Soup w/ Average 51.8 / 66.4 38.8 / 51.7 28.0 / 39.4
Encoder-Decoder Finetuned 8B
Full Finetuned - Average Pooling With & Without Norms
Soup w/ Average 55.8 / 69.8 47.8 / 61.3 38.7 / 50.9
Average + Norm Before 35.9 / 47.8 47.8 / 60.9 38.1 / 50.2
Average + Norm After 50.4 / 65.2 42.3 / 55.6 33.4 / 44.9
Average + Norm Before & After 6.9 / 10.7 42.2 / 54.7 33.8 / 45.6
Full Finetuned - Summation Pooling With & Without Norms
Soup w/ Sum 55.1 / 69.4 44.2 / 57.4 25.0 / 36.0
Sum + Norm Before 8.8 / 13.6 8.1 / 13.0 4.6 / 9.4
Sum + Norm After 51.9 / 66.1 43.6 / 56.6 34.7 / 46.2
Sum + Norm Before & After 10.2 / 16.3 40.3 / 52.9 33.2 / 44.8
Full Finetuned - Max Pooling Without Norms
Soup w/ Max 52.3 / 65.6 39.1 / 51.6 28.9 / 40.5
Table 1: HotpotQA performance (Exact Match / F1) for Mamba2-8B models trained on 5 documents (2 gold + 3 distractors) and evaluated on $n$ documents, each with 2 gold and $(n{-}2)$ distractors. We compare pretrained models, decoder-only finetuning, and encoder-decoder finetuning across different pooling strategies (average, sum, max), with optional normalization applied before and/or after aggregation. Underlined entries indicate evaluations where the number of test-time documents matches the training configuration. We observe that decoder-only finetuning improves performance by learning to interpret fixed souped states, and full encoder-decoder finetuning yields the best results by also learning to produce mergeable representations. Across all configurations, simple averaging without normalization emerges as the most stable and effective aggregation method.
3 Experiments
We evaluate state souping across a range of long-context reasoning tasks that test different dimensions of soupability: single-hop vs. multi-hop inference, monolithic vs. multi-document structure, and sparse signal detection in distractor-heavy inputs.
We study multi-document question answering in two distinct regimes. For single-hop QA, we use the QA subset of the RULER dataset [11], which augments SQuAD-style [19] questions for long-context evaluation. Each question is answerable from a single gold document placed within a large set of distractors, allowing us to isolate how well the model can identify and preserve localized information in souped representations. For multi-hop QA, we turn to HotpotQA [28], a benchmark requiring compositional reasoning across multiple Wikipedia paragraphs. Each question demands integration of evidence from at least two documents, providing a direct test of whether souped hidden states preserve the relational structure needed for multi-hop reasoning.
We also evaluate on long-document QA tasks, where inputs are single extended narratives rather than disjoint documents. In this setting, we train on the RACE dataset [14], which consists of relatively short educational passages, and test on both RACE and the validation portion of QuALITY [16], which features much longer and more complex narratives. This setup allows us to assess whether document segmentation and aggregation via souping generalizes from short-form to long-form content. Our findings support earlier observations from the QuALITY paper that models trained on RACE can transfer effectively.
Finally, we evaluate sparse retrieval using the multikey-2 subset of the RULER dataset. In this setting, each document line contains a unique identifier paired with a corresponding value, and the model is tasked with memorizing all such mappings. At inference time, it receives a query specifying one of the identifiers and must output the correct value. The full input includes many such mappings, requiring the model to store a dense set of key-value pairs and retrieve the correct one after aggregation. This task challenges the model’s ability to preserve fine-grained, instance-specific information across multiple independently encoded documents.
Together, these tasks span diverse input formats and reasoning types, enabling a broad and controlled evaluation of state souping. Unless otherwise noted, all experiments are conducted using Mamba2-8B. Results for the 2.7B model are reported in Appendix B. In general, we observe that the larger model consistently achieves stronger performance, particularly in settings requiring multi-hop reasoning or generalization across segment length.
Table 2: HotpotQA results for Mamba2-8B evaluated on 2 gold documents plus $(n{-}2)$ distractors. We compare concat-based and soup-based training across a range of finetuning configurations: QA-only (no documents), 2-gold only, and 2-gold + distractors. Each model is evaluated at multiple soup sizes, and underlined entries mark evaluations that match the training number of documents.
Method Test on $2$ gold + $(n{-}2)$ distractors
$n=0$ 2 3 4 5 6 7 8 9 10
Pretrained, no finetune
Concat 2.4 / 4.7 15.4 / 26.3 11.1 / 22.5 9.7 / 21.6 8.5 / 20.2 7.2 / 18.9 6.4 / 17.1 5.5 / 16.5 6.0 / 17.1 5.0 / 15.7
[…]
Full Finetuned on $n=0\ gold+0\ distractors$ (QA-only)
Concat 18.8 / 27.4 52.2 / 67.4 48.0 / 62.3 44.9 / 58.5 42.0 / 55.1 38.5 / 51.5 36.6 / 49.5 36.3 / 48.8 34.8 / 46.9 33.0 / 45.0
Full Finetuned on $n=2\ gold+0\ distractors$ documents
Concat 56.0 / 70.3 51.3 / 65.1 46.2 / 59.8 43.6 / 57.3 40.1 / 53.5 37.6 / 50.5 36.2 / 48.1 34.4 / 46.4 34.4 / 45.8
[…]
Full Finetuned on $n=2\ gold+3\ distractors$ documents
Concat 57.1 / 71.3 54.4 / 68.3 51.9 / 66.3 49.0 / 63.1 47.9 / 61.5 45.5 / 59.0 44.4 / 57.8 42.9 / 55.4 41.4 / 54.1
[…]
Full Finetuned on $n=2\ gold+5\ distractors$ documents
Concat 54.6 / 69.4 52.6 / 67.1 50.8 / 64.4 49.1 / 63.1 47.8 / 61.4 45.1 / 58.7 44.3 / 57.3 42.4 / 55.8 41.6 / 54.3
[…]
Full Finetuned on $n=2\ gold+8\ distractors$ documents
Concat 55.0 / 69.5 52.1 / 66.7 48.9 / 63.5 48.0 / 62.5 46.9 / 60.8 45.2 / 59.0 43.2 / 57.6 42.0 / 56.0 42.5 / 56.0
[…]
$n=0$ corresponds to a no-context setting where the model receives only the query (i.e., 0 gold documents and 0 distractors).
Table 3: EM (%) on the NIAH task for Mamba-2 8B models finetuned on 25K examples with either 4K (left) or 8K (right) sequence length. Models are evaluated across varying numbers of document segments and sequence lengths. Bold indicates the best score in each column. Gray cells mark results outperforming the respective concat-finetuned baseline (85.8 / 24.25 for 4K, 81.65 / 35.55 for 8K).
Finetuned on 4K sequence length
Method Train Segments Test Segments Test Seq. Length
4k 8k
Concat 1 1 85.8 24.25
Soup w/ Average 2 2 87.0 38.45
4 79.75 32.05
8 45.7 16.3
16 2.3 0.8
32 0.0 0.0
[…]
4 86.8 38.7
8 76.55 29.3
16 29.05 11.45
32 0.9 0.5
[…]
4 84.9 36.85
8 83.6 36.45
16 71.45 28.9
32 15.75 9.25
Finetuned on 8K sequence length
Method Train Segments Test Segments Test Seq. Length
4k 8k
Concat 1 1 81.65 35.55
Soup w/ Average 2 2 89.8 48.4
4 79.45 38.15
8 41.15 15.45
16 3.2 1.45
32 0.05 0.0
[…]
4 86.5 43.85
8 75.65 34.6
16 23.8 11.5
32 0.4 0.55
[…]
4 88.2 45.85
8 85.25 43.9
16 64.95 32.25
32 9.7 7.9
3.1 Experimental Setup
3.1.1 Training Configurations
To ensure reproducibility, we fix all random seeds to 42 and enable deterministic=True. Optimization follows the setup from [8], using AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.95$, gradient clipping at 1.0, and a cosine decay schedule with 10% linear warm-up. The 2.7B model is finetuned with a learning rate of $5\times 10^{-5}$, and the 8B model with $1\times 10^{-5}$. All experiments are trained for a single epoch, with one exception: the single-hop QA setting. For single-hop QA, we constructed a dataset of approximately 6,000 examples by augmenting SQuAD-style questions using the RULER framework. Of these, 5,000 examples were used for training. Due to the smaller dataset size, one epoch of training was insufficient for convergence, so we trained for 3 epochs in this setting.
Gradient accumulation is set to 120 steps111128 for the 2.7B HotpotQA configuration.. For distributed training, we use DeepSpeed ZeRO-2 [18], which allows each GPU to hold the full model and facilitates the extraction of complete hidden states for the souping mechanism. Experiments were mainly conducted on clusters with H100 and H200 GPUs; Full training and implementation details are provided in Appendix A.
3.1.2 Evaluation Metrics
We evaluate model performance using standard metrics for both extractive and multiple-choice question answering. For extractive QA tasks such as HotpotQA and RULER, we report exact match (EM), which measures the percentage of predictions that exactly match any ground-truth answer span, and F1 score, which reflects the token-level overlap between the predicted and reference answers. The F1 score is defined as the harmonic mean of precision and recall:
[…]
For multiple-choice QA tasks, including RACE and QuALITY, we use multiple-choice accuracy as the evaluation metric. In this setting, the model outputs a distribution over the four answer options (A, B, C, D), and we compute the log-probability of each choice. The predicted answer corresponds to the option with the highest score, and accuracy is measured as the proportion of correctly selected answers.
3.1.3 Baselines and Reference Setting
We compare souping against several baselines that represent different levels of supervision and context integration. First, we include a pretrained baseline: an off-the-shelf Mamba2 model used without any task-specific finetuning. This setup serves as a lower bound, as the model has not been adapted to the QA tasks.
Next, we consider a QA-only finetuning baseline. Here, the model is trained only on $(q,a)$ pairs without access to any document context during training. Although this configuration does not support long-context integration, it provides a useful reference for how much performance is attributable to the context.
Finally, we evaluate a concat-based finetuning baseline, where the model is trained end-to-end on full input sequences—i.e., $(d_{1},\dots,d_{k},q,a)$—concatenated into a single flat input. This setting provides strong supervision by exposing the model to all relevant documents in a joint context window and serves as our primary reference point when assessing the performance of souping-based alternatives.
Our aim is to demonstrate that, with appropriate finetuning, souping-based models can outperform context-free baselines and approach the performance of the concat-based reference model.
Table 4: Multiple-choice QA accuracy (%) for Mamba-2 8B finetuned on 25K RACE examples. Test is performed on RACE (5K) and QuALITY (2K) using answer selection based on maximum choice probability. Bolded values denote the highest score in each test. Gray cells indicate cases where soup-based finetuning outperforms concat-based finetuning.
Method Train Segments Test Segments on RACE (5K) Test Segments on Quality (2K)
0 1 2 4 8 16 0 1 2 4 8 16
Concat (QA-only) 0 26.54 26.08
Concat (With Context) 1 56.01 32.69
Soup w/ Average 2 63.62 59.43 56.49 52.30 47.46 47.27 45.16 43.10
4 66.02 63.54 60.58 57.31 51.15 50.05 48.08 45.69
8 62.29 59.28 55.74 53.77 46.31 45.69 44.20 43.29
3.2 Training Strategies for Soupability
We evaluate three training regimes that differ in how the model is exposed to and trained on soupable hidden states.
In the out-of-the-box setting, we merge hidden states from a pretrained Mamba2 encoder and pass them to the decoder without any finetuning. As expected, performance is poor, as the model has not learned to interpret pooled representations.
In the decoder-only finetuning setup, the encoder is frozen and only the decoder is trained to answer questions from fixed, aggregated states. This tests whether the decoder alone can learn to interpret pooled inputs. Performance improves substantially over the baseline but remains constrained by the unadapted encoder outputs.
Full encoder-decoder finetuning trains both components jointly, allowing the encoder to produce more mergeable states and the decoder to interpret them effectively. As shown in Table 1, this setup consistently achieves the strongest results, confirming that SSMs can be trained end-to-end to support modular reasoning through state pooling.
3.3 Aggregation Strategy Comparison
We evaluate several strategies for aggregating per-layer hidden states across documents, including elementwise average, sum, and maximum, along with optional unit normalization applied before or after aggregation. Results are reported in Table 1.
When evaluating a model trained on 5 documents and tested across varying soup sizes, we observe that overall performance is relatively stable across aggregation methods, and no-norm averaging consistently yields strong performance. This robustness suggests that averaging provides a natural balance: it aggregates across documents while keeping the scale of the hidden states stable.
By contrast, summation without normalization tends to degrade more quickly as more documents are merged. We hypothesize that this is due to the unbounded growth in activation magnitude. Applying normalization after summation improves performance at larger document counts. However, it still underperforms averaging, which implicitly maintains consistent magnitude while preserving directional information. Interestingly, normalization before aggregation performs well with averaging but poorly with summation, likely because it destroys relative magnitude differences across documents that are helpful when using sum.
Taken together, these observations suggest that averaging without normalization offers the best trade-off between simplicity, stability, and generalization. We adopt this as the default aggregation method in all subsequent experiments.
Table 5: EM / F1 scores on RULER QA_1 for Mamba2-8B trained and evaluated on 4k sequence length. Gray cells indicate performance exceeding the concat-finetuned test results (EM 54.81 / F1 71.90), and bold marks the highest test result of the task. Training on more segments improves generalization to higher evaluation segment counts.
Method Train Test Segments
Segments 1 2 5 10 20
Concat 1 54.81 / 71.90
Soup w/ Average 2 58.38 / 74.05 34.89 / 49.04 13.19 / 23.40 12.36 / 22.45
5 58.38 / 73.89 57.28 / 72.09 35.71 / 50.80 14.97 / 26.25
10 21.02 / 31.41 13.87 / 22.36 52.75 / 68.68 43.82 / 58.99
20 28.71 / 41.44 28.85 / 42.93 45.60 / 61.78 44.37 / 60.85
3.4 Generalization Across Soup Sizes
We study how the number of documents or segments being merged at inference time affects model performance, particularly when this number differs from the training configuration. This analysis addresses the extent to which soupability generalizes across varying soup sizes.
Tables 2, 3, 4, and 5 show that performance typically peaks when the number of merged inputs at inference matches the training-time setting. Increasing the soup size beyond the training configuration leads to a gradual performance drop, likely due to information dilution or representational interference. Reducing the number of inputs typically causes a smaller degradation, suggesting some robustness to under-composition. Notably, models trained with larger soup sizes exhibit better generalization to even larger settings. For example, a model trained on 8 segments maintains strong performance when evaluated on 16 segments, whereas a model trained on only 2 segments generalizes poorly in this setting. This trend is consistent across both QA and retrieval tasks.
These results suggest that exposure to more fragmented input during training helps the model learn representations that are more tolerant to variation in soup size. When paired with appropriate finetuning, this property enables robust reasoning over multi-document contexts of varying width.
3.5 Effect of Corpus Structure on Soupability
We evaluate how different input structures—specifically long single-document narratives and multi-document corpora—affect the effectiveness of state souping.
Long Documents.
To assess whether souping is viable for long, contiguous inputs, we experiment with the RACE and QuALITY datasets. RACE consists of relatively short reading passages, while QuALITY includes significantly longer and more complex narratives. Both are evaluated in a multiple-choice QA setting.
Interestingly, models trained with souping on RACE generalize well to QuALITY (Table 4), despite the large increase in input length. This is likely due to the structure of SSMs: each document’s hidden state is a fixed-size vector, independent of the input length. As a result, the model can continue to aggregate representations effectively, even as document or segment length increases. In contrast, models finetuned using standard concatenation on RACE generalize poorly to QuALITY, where the input sequence length grows beyond what the model was exposed to during training. This suggests that souping offers a more length-invariant mechanism for representation and generalization.
Multi-hop Reasoning.
We also examine whether souping supports compositional reasoning over multiple documents, as required by HotpotQA. As shown in Table 2, models trained with souping can perform multi-hop reasoning across independently encoded documents, achieving performance on par with concat-based models. This indicates that the merged hidden states preserve cross-document relationships without requiring joint encoding.
Together, these results highlight that souping is robust to input granularity—from short segments to long narratives—and supports both single-hop and multi-hop QA. This flexibility is critical for applications that must operate over heterogeneous corpora without the cost of re-encoding or sequence length tuning.
4 Related Work
Long-Context Sequence Modeling.
Transformers [21] remain the dominant architecture for modeling long-range dependencies, but their $\mathcal{O}(L^{2})$ attention cost and $\mathcal{O}(L)$ memory growth in the KV cache hinder scalability. A rich body of work explores architectural modifications to overcome this, including sparse attention [4, 1, 29], linearized attention [12], and chunked processing with memory compression [5, 26, 27]. Retrieval-augmented generation methods such as RAG [15] provide complementary approaches by offloading long-context memory to external sources.
Structured State Space Models (SSMs) offer an alternative path with fundamentally different scaling characteristics. Models like S4 [10, 9] introduced linear-time computation via structured recurrences. Mamba [8] builds on this with input-dependent gating, while Mamba2 [6] unifies SSMs and attention through structured state space duality, yielding substantial speedups on long sequences with strong downstream performance. Recent work further demonstrates the utility of Mamba-based models for dense and retrieval-augmented tasks, including dense passage ranking [31] and long-document retrieval in RAG pipelines [3].
Model and Task Souping.
Model souping refers to merging parameters across finetuned checkpoints to improve robustness and generalization without retraining [25, 20]. Recent work has explored souping internal representations rather than weights. State Soup [17] linearly combines task-specific hidden states for skill transfer. In contrast, our method focuses on merging hidden states from disjoint document chunks—enabling compositional reasoning across distributed corpora through simple aggregation strategies.
Parallel and Distributed Ingestion.
Efficient ingestion of long contexts has been addressed through sparse or structured attention mechanisms [4, 29], memory compression across chunks [5, 26], and retrieval-based pipelines [15]. For SSMs, constant-memory recurrence and linear compute make them naturally suited to chunk-wise streaming. However, to our knowledge, no prior work has explicitly studied post-hoc merging of hidden states across chunks for joint reasoning.
RNN-Inspired State Mixing.
Recurrent models have long supported state-passing across time steps, and early works explored implicit mixing of information through recurrent transitions [7]. Our approach differs in that it investigates explicit aggregation of intermediate hidden states produced in parallel, rather than temporal chaining of recurrent updates.
Key/Value Cache Merging.
Recent work has explored adaptive merging of key/value caches to compress context without retraining [24], and memory-efficient cache eviction policies [32]. While related in spirit, these approaches focus on attention-based caches, whereas our work targets SSM hidden states.
Token Merging in Vision Transformers.
In the vision domain, token merging has been studied as a means of compressing long sequences. TCFormer [30] clusters tokens across space for progressive downsampling. Other work merges tokens dynamically based on content [2, 13], or selects value/key pairs to retain or evict during inference [23]. These ideas echo our motivation to aggregate compact, composable representations—though in our case applied to document states within state space models.
5 Conclusion
We present document souping, a method for merging hidden states across independently encoded documents in structured state space models. By aggregating layer-wise representations through simple commutative operations, souping enables modular, parallel ingestion of long-context inputs while supporting accurate downstream inference. Our experiments demonstrate that Mamba2 models trained with souping are capable of multi-hop reasoning, sparse retrieval, and generalization across document and segment scales—often matching or outperforming traditional concat-based finetuning.
While the results are promising, our study has several limitations. First, we were unable to evaluate soupability at very large segment counts (e.g. 128) due to our compute and memory constraints; investigating the limits of scalability in high-fragmentation regimes remains an open question. Second, we did not benchmark training or inference efficiency directly. Such analysis would require large-scale deployment infrastructure to meaningfully measure throughput and memory usage under realistic workloads, which we leave to future work.
Overall, our findings highlight state souping as a lightweight and effective strategy for long-context reasoning in SSMs, and we hope it provides a foundation for broader adoption and further exploration.
Appendix A Implementation Details
A.1 Data Formatting
To support both standard fine-tuning and our proposed souping method, we define two input formats:
  • Concat-data: the input $x$ is formed by linearly concatenating $k$ documents, the question $q$, the answer $a$, and an end-of-sequence token $\langle\mathrm{eos}\rangle$:
    $x=d_{1}\,\mathbin{\oplus}\,d_{2}\,\mathbin{\oplus}\,\cdots\,\mathbin{\oplus}\,d_{k}\,\mathbin{\oplus}\,q\,\mathbin{\oplus}\,a\,\mathbin{\oplus}\,\langle\mathrm{eos}\rangle.$
    This aligns with conventional LM fine-tuning, where the model attends jointly over the entire sequence.
  • Souping-data: identical to concat-data, except that after each document $d_{i}$ we insert a special separator token $\langle\mathrm{DOC\_SEP}\rangle$. Formally,
    $x=d_{1}\,\mathbin{\oplus}\,\langle\mathrm{DOC\_SEP}\rangle\,\mathbin{\oplus}\,d_{2}\,\mathbin{\oplus}\,\langle\mathrm{DOC\_SEP}\rangle\,\mathbin{\oplus}\,\cdots\,\mathbin{\oplus}\,d_{k}\,\mathbin{\oplus}\,\langle\mathrm{DOC\_SEP}\rangle\,\mathbin{\oplus}\,q\,\mathbin{\oplus}\,a\,\mathbin{\oplus}\,\langle\mathrm{eos}\rangle.$
    During preprocessing, we split on $\langle\mathrm{DOC\_SEP}\rangle$ to isolate each document chunk for parallel encoding.
Notation.
[…]
Algorithm 1 Input Formatting for Concat- vs. Souping-data
1:Documents $\{d_{1},\ldots,d_{k}\}$, question $q$, answer $a$, flag $\mathrm{soup}\in\{\mathrm{true},\mathrm{false}\}$
2:Formatted input $x$
3:$x\leftarrow$ empty string
4:for $i=1$ to $k$ do
5:     if $\mathrm{soup}$ then
6:         $x\leftarrow x\mathbin{\oplus}d_{i}\mathbin{\oplus}\langle\mathrm{DOC\_SEP}\rangle$
7:     else
8:         $x\leftarrow x\mathbin{\oplus}d_{i}\mathbin{\oplus}\texttt{" "}$
9:     end if
10:end for
11:$x\leftarrow x\mathbin{\oplus}q\mathbin{\oplus}\texttt{" "}\mathbin{\oplus}a\mathbin{\oplus}\langle\mathrm{eos}\rangle$
A.2 Souping Recipe
A.2.1 Without Activation Checkpointing
Given a souping-data sequence $x$, we split at each $\langle\mathrm{DOC\_SEP}\rangle$, producing $k+1$ segments:
$\{s_{1},\dots,s_{k},s_{k+1}\},$
where $s_{i}=d_{i}$ for $i\leq k$ and $s_{k+1}$ contains $(q,a,\langle\mathrm{eos}\rangle)$. We then form:
  1. 1.
    Doc-batch formation. Stack the $k$ context segments $\{s_{1},\dots,s_{k}\}$ along the batch dimension, yielding a tensor of shape $(B\times k,T,D)$, where $B$ is the original mini-batch size, $T$ is the (left-padded) segment length, and D is embedding dimension.
  2. 2.
    Parallel encoding. Pass the Doc-batch through the encoder in parallel to obtain per-document hidden states $\{h_{i}\}_{i=1}^{B\times k}$ for each batch element.
  3. 3.
    Aggregation. Merge documents $\{h_{i}\}$ for each sequence of the mini-batch via the chosen souping strategy (average, sum, or max) to produce a single souped state $\{h_{\mathrm{soup\_j}}\}_{j=1}^{B}$ per batch element.
  4. 4.
    QA-batch processing. Take the QA segment $s_{k+1}$, right-pad it to length $T$ to form a QA-batch of shape $(B,T,D)$. We feed the QA-batch along with $h_{\mathrm{soup}}$ injected at every layer—through the decoder. The souped state remains constant until decoding begins.
  5. 5.
    Loss computation. Map decoder outputs back to token positions, convert to logits, and compute cross-entropy loss over the answer tokens.
A.2.2 With Activation Checkpointing
To support training with more documents and larger model sizes without running out of memory, we employ document-level activation checkpointing. Unlike the parallel encoding scheme in Section A.2.1, which processes all context documents of a mini-batch simultaneously, we encode one document per mini-batch at a time under checkpointing.
Specifically, we iterate over the $k$ context segments for each batch and perform a forward pass for a single document per step, using PyTorch’s gradient checkpointing to trade off compute for memory. During backpropagation, each document’s encoder activations are recomputed on-the-fly. This significantly reduces peak memory usage, enabling us to scale to more documents (higher $k$), longer input sequences, and larger model weights.
While this sequential encoding incurs additional compute overhead, it offers a practical trade-off to unlock training regimes otherwise inaccessible due to memory constraints. If sufficient GPU memory is available (e.g., large HBM capacity), the overhead of activation checkpointing can be amortized by increasing the batch size—sometimes even resulting in faster end-to-end training than without checkpointing, due to improved throughput and parallelism.
A.3 Pseudo-code
Algorithm 2 describes the inference-time procedure for corpus encoding via document souping. Each document $d_{i}$ is processed independently by a shared SSM encoder to produce a set of per-layer hidden states $\{h_{i}^{(1)},\dots,h_{i}^{(L)}\}$. Optionally, each hidden state can be normalized (e.g., to unit norm) before being appended to a layer-specific collection. After all documents are encoded, hidden states are aggregated across documents at each layer using a commutative pooling operation such as averaging, summation, or max. The resulting pooled states $\{h_{\text{soup}}^{(1)},\dots,h_{\text{soup}}^{(L)}\}$ form a compact, merged representation of the document set, which is then used to condition the decoder when answering a query $q$. The document encoding can be performed offline and cached in advance, enabling efficient reuse across queries.
Algorithm 2 Corpus Encoding via Document Souping (Inference Only)
1:Document set $\{d_{1},\dots,d_{k}\}$, query $q$
2:Predicted answer $\widehat{y}$
3:Initialize empty list of layerwise states $\{H^{(l)}\}_{l=1}^{L}$
4:for each document $d_{i}$ do
5:     Compute hidden states: $\{h_{i}^{(1)},\dots,h_{i}^{(L)}\}\leftarrow\mathrm{SSM}_{\theta}(d_{i})$
6:     for $l=1$ to $L$ do
7:         Optionally normalize: $\tilde{h}_{i}^{(l)}\leftarrow h_{i}^{(l)}/\|h_{i}^{(l)}\|$
8:         Append $\tilde{h}_{i}^{(l)}$ to $H^{(l)}$
9:     end for
10:end for
11:for $l=1$ to $L$ do
12:     Pool document states: $h_{\text{soup}}^{(l)}\leftarrow\mathrm{pool}(H^{(l)})$
13:     Optionally normalize: $h_{soup}^{(l)}\leftarrow h_{soup}^{(l)}/\|h_{soup}^{(l)}\|$
14:end for
15:Predict: $\widehat{y}\leftarrow\mathrm{SSM}_{\theta}\left(q\mid\{h_{\text{soup}}^{(l)}\}_{l=1}^{L}\right)$
A.4 Decoding Hyperparameters
[…]
Soup achieved best validation performance with temperature $=0.3$, top-$p=0.5$, and top-$k=20$, while Concat performed optimally with temperature $=0.5$, top-$p=0.95$, and top-$k=30$. All subsequent experiments adopted these respective configurations for each method.
Table 6: HotpotQA results for Mamba-2 2.7B with $2$ gold documents and $n-2$ distractors. Cells show EM / F1. Underline marks models tested on the same docs they are trained on.
Model Evaluation: $2$ gold + $(n{-}2)$ distractors
$n=0$ 2 3 4 5 6 7 8 9 10
Pretrained, no finetune
Concat 3.3 / 7.8 8.4 / 22.7 5.8 / 18.6 5.2 / 17.1 3.5 / 14.0 3.2 / 12.9 2.8 / 11.8 2.2 / 11.6 2.2 / 11.0 2.3 / 10.4
Finetuned on $n=0\ gold+0\ distractors$ (QA-only)
Concat 12.7 / 19.4 40.9 / 54.2 36.7 / 49.1 32.4 / 44.7 30.4 / 41.8 28.0 / 39.2 26.2 / 36.9 25.4 / 35.5 23.6 / 33.7 22.3 / 32.2
Finetuned on $n=2\ gold+0\ distractors$ documents
Concat 50.8 / 65.2 46.0 / 59.2 41.7 / 54.5 37.7 / 49.7 35.9 / 47.5 33.8 / 44.6 30.1 / 40.9 28.4 / 38.4 26.6 / 36.9
[…]
Finetuned on $n=2\ gold+3\ distractors$ documents
Concat 48.8 / 63.5 46.3 / 60.2 43.1 / 56.8 42.3 / 55.2 40.6 / 53.6 37.1 / 49.6 35.3 / 48.0 35.2 / 47.6 33.7 / 45.6
[…]
Finetuned on $n=2\ gold+5\ distractors$ documents
Concat 50.4 / 64.4 47.0 / 60.9 42.7 / 56.4 41.8 / 55.0 40.1 / 52.8 38.1 / 50.3 37.1 / 49.4 36.9 / 48.8 34.8 / 46.7
[…]
Finetuned on $n=2\ gold+8\ distractors$ documents
Concat 49.7 / 63.7 45.2 / 58.8 43.6 / 56.7 41.7 / 54.1 40.0 / 53.1 36.8 / 48.8 36.4 / 48.4 35.3 / 47.4 33.9 / 45.7
[…]
$n=0$ corresponds to a no-context setting where the model receives only the query (i.e., 0 gold documents and 0 distractors).
A.4.1 Compute Resources
Experiments were conducted using a combination of on-premise and cloud nodes, equipped with high-memory GPUs and multi-core CPUs. All runs used PyTorch 2.6 with CUDA 12.4.
  • Node A (On-premise): 8× NVIDIA L40S (48GB) with AMD EPYC 7282 CPU
  • Node B (On-premise): 8× NVIDIA H100 (80GB) with Intel Xeon Platinum 8480+ CPU
  • Node C (Cloud): On-demand NVIDIA H200 (144GB) with AMD EPYC 9654 CPU
Nodes A and B were used for model development, ablations, and full training runs. Node C enabled rapid parallel experimentation.
Training time varies with model size, input sequence length, number of documents per example, and total steps. For instance, fine-tuning an 8B model on HotpotQA (30K examples, 5 documents per example) requires approximately 3 hours on a single H200 GPU with activation checkpointing enabled.
Multi-document souping increases memory overhead. The following configurations represent minimum requirements for training on HotpotQA with 5-document inputs and small batch sizes:
  • 2.7B model: $\geq$ 1× 80GB or 2× 48GB GPUs
  • 8B model: $\geq$ 1× 144GB, 2× 80GB, or 4× 48GB GPUs
Training with longer input sequences or larger segments proportionally increases memory demands, requiring additional GPUs or higher-capacity HBM.
Appendix B Additional Results and Analysis
B.1 HotPotQA Extended Analysis
B.1.1 Mamba2-2.7B Results
Table 6 shows HotpotQA results for Mamba2-2.7B across a range of training and evaluation soup sizes. We observe that performance improves consistently with fine-tuning and scales with the number of documents seen during training. However, compared to the 8B model (Table 2), the 2.7B model shows greater sensitivity to soup size mismatch and a more pronounced performance drop at larger test-time document counts. This supports our finding that larger models generalize more robustly across soup sizes and better tolerate distribution shift during inference.
Figure 2: Exact Match (EM) scores on HotpotQA for Mamba2-8B evaluated across increasing numbers of input documents. Each line represents a model trained on 5 documents (2 gold + 3 distractors) with a different pooling and normalization configuration. Soup w/ Average consistently remains robust across all tested document sizes compared to other configurations.
Figure 3: F1 scores on HotpotQA for Mamba2-8B using the same experimental setup as Figure 2.
B.1.2 Mamba2-8B Results
Table 7 expands on the main paper’s Table 1 by providing a more detailed view of HotpotQA results for the Mamba2-8B model trained on 5 documents (2 gold + 3 distractors). At inference time, we evaluate the same checkpoint across a range of input sizes, including a no-context setting (query-only) and settings with 2 to 10 documents, where each includes 2 gold documents and $n{-}2$ distractors.
The table includes comparisons across several pooling operators (average, sum, max) and normalization variants, and includes pretrained, decoder-only, and full encoder-decoder fine-tuning settings. The results show that while all models degrade slightly as the number of input documents increases, full encoder-decoder fine-tuning with no-norm averaging consistently yields the best performance across soup sizes. Overall, these results highlight the robustness of soup-based representations under increasing input width and reinforce the conclusions drawn in the main text.
To visualize these trends, Figure 2 and Figure 3 plot EM and F1 scores, respectively, across different input sizes. These plots confirm that while pretrained and decoder-only models degrade rapidly, full finetuning with average pooling remains more robust. The visualizations emphasize how different configurations respond to increasing distractor load, further illustrating the stability of soup representations under input variation.
Table 7: HotpotQA performance (Exact Match / F1) for Mamba-2 8B trained on 5 documents (2 gold + 3 distractors) and tested on $n$ documents, with $2$ gold and $(n-2)$ distractors. We compare pretrained models, decoder-only fine-tuning, and encoder-decoder fine-tuning, under different pooling operators (Avg, Sum, Max) and normalization settings. Underline marks models tested on the same docs they are trained on.
Method Test on $2$ gold + $(n-2)$ distractors
2 3 4 5 6 7 8 9 10
Pretrained 8B (No Finetune)
Concat 15.4 / 26.3 11.1 / 22.5 9.7 / 21.6 8.5 / 20.2 7.2 / 18.9 6.4 / 17.1 5.5 / 16.4 6.0 / 17.1 5.0 / 15.7
[…]
Decoder-Only Finetuned 8B
Soup w/ Average 51.8 / 66.4 46.7 / 60.1 42.2 / 55.6 38.8 / 51.7 35.2 / 48.1 33.5 / 46.0 31.4 / 43.7 29.6 / 41.3 28.0 / 39.4
Encoder-Decoder Finetuned 8B
Full Finetuned - Average Pooling With & Without Norms
Soup w/ Average 55.8 / 69.8 53.0 / 66.5 50.0 / 63.6 47.8 / 61.3 45.3 / 58.7 43.9 / 56.6 40.9 / 53.6 39.5 / 52.2 38.7 / 50.9
[…]
Full Finetuned - Summation Pooling With & Without Norms
Soup w/ Sum 55.1 / 69.4 51.4 / 65.3 47.2 / 61.2 44.2 / 57.4 40.3 / 53.3 37.0 / 49.4 34.1 / 46.2 29.2 / 40.9 25.0 / 36.0
[…]
Full Finetuned - Max Pooling Without Norms
Soup w/ Max 52.3 / 65.6 46.4 / 59.6 42.4 / 55.1 39.1 / 51.6 36.6 / 48.6 34.1 / 46.1 32.6 / 44.6 31.1 / 42.4 28.9 / 40.5
B.2 Needle in a Haystack
Table 8 shows NIAH accuracy for Mamba2-2.7B Finetuned on either 4k or 8k sequence lengths using 25K examples. Results are shown across a grid of train/test segment combinations. Models trained with more segments generalize better to larger soup sizes during inference, with improvements especially pronounced at higher segment counts. In contrast, models trained on just 2 segments degrade sharply when evaluated on 8, 16, or 32 segments, highlighting the importance of training with sufficient segments for robust generalization in soup-based Finetuned models.
Table 8: Accuracy on EM (%) on the NIAH task for Mamba2-2.7B models trained with 4K (left) and 8K (right) sequence lengths for 25K examples. Bold indicates the best result in each column. Gray cells indicate accuracy exceeding the Concat-finetuned baseline (84.4 / 36.1 for 4K, 78.3 / 43.05 for 8K).
Finetuned on 4K Sequence Length
Method Train Segments Test Segments Test Seq. Length
4k 8k
Concat 1 1 84.4 36.1
Soup w/ Avg. 2 84.4 33.9
4 71.7 35.5
2 8 26.2 12.9
16 0.3 0.3
32 0.0 0.0
[…]
4 86.5 49.2
4 8 72.1 39.2
16 25.2 13.1
32 0.2 0.4
[…]
4 82.8 39.9
8 8 81.4 43.9
16 66.5 32.7
32 20.2 9.9
[…]
4 77.1 30.6
16 8 83.5 40.5
16 82.4 45.3
32 70.2 35.8
Finetuned on 8K Sequence Length
Method Train Segments Test Segments Test Seq. Length
4k 8k
Concat 1 1 78.3 43.05
Soup w/ Avg. 2 84.4 59.2
4 71.6 47.8
2 8 28.5 15.1
16 0.2 0.3
32 0.0 0.0
[…]
4 82.5 59.3
4 8 64.7 42.5
16 13.5 10.0
32 0.1 0.2
[…]
4 85.3 59.2
8 8 80.7 57.1
16 60.3 41.0
32 9.7 7.5
[…]
4 86.6 54.6
16 8 86.2 60.3
16 81.0 59.0
32 58.7 42.1
Table 9 reports NIAH accuracy for Mamba2-8B trained with only 7K examples at either 4k or 8k sequence length. Despite the reduced supervision and limited training (1 epoch), the 8B model achieves strong results and continues to exhibit the same trends as in higher-resource settings: soup-finetuned models outperform concat-finetuned baselines, and training with more segments improves generalization to higher test-time segment counts. In contrast, we found that the Mamba2-2.7B model did not converge reliably under this low-data setup. These results suggest that soupability and generalization benefits persist even in lower-resource regimes, but larger models are more robust to limited training data.
Table 9: Evaluation accuracy (%) on NIAH task for 8B Mamba models Finetuned on 4k (left) or 8k (right) sequence length with 7K examples. Gray means Soup-finetuned results are better than respective Concat-finetuned results (72.55 / 14.2 for 4k, 66.55 / 18.7 for 8k). Bold represents the best in each column.
Finetuned on 4K sequence length
Method Train Segments Test Segments Test Seq. Length
4k 8k
Concat 1 1 72.55 14.2
Soup w/ Avg. 2 2 80.4 30.2
4 70.55 23.7
8 38.0 10.65
[…]
4 75.4 25.35
8 58.6 18.5
[…]
4 73.5 24.55
8 68.8 22.35
Finetuned on 8K sequence length
Method Train Segments Test Segments Test Seq. Length
4k 8k
Concat 1 1 66.55 18.7
Soup w/ Avg. 2 2 74.15 31.0
4 57.75 21.55
8 18.85 7.5
[…]
4 69.7 27.95
8 48.8 17.85
[…]
4 73.0 27.35
8 67.1 25.05
B.3 Ruler QA_1
We trained for 3 epochs and tested Mamba2-2.7B on the QA_1 subset of RULER, where each input includes 20 documents, only one of which contains the answer. This single-hop task tests the model’s ability to isolate relevant information under extreme distraction. As shown in Table 10, Soup w/ Average trained on 2 segments slightly outperforms the concat baseline. While the margin is small and not statistically significant, this result demonstrates the potential of state souping even in sparse-relevance QA settings.
However, training on larger segment counts leads to worse performance, especially on smaller test-time soup sizes. Unlike multi-hop settings like HotpotQA, where broader exposure improves generalization, over-fragmentation in sparse single-hop tasks appears to dilute the signal.
Table 10: EM / F1 scores on RULER QA_1 for Mamba2-2.7B trained and evaluated on 4k sequence length. Gray cells indicate performance exceeding the concat-finetuned test results (EM 48.4 / F1 64.8), and bold marks the highest test result of the task. Training on more segments improves generalization to higher evaluation segment counts until train with 20 segments. All experiments are run for 3 epochs due to limited data.
Method Train Test Segments
Segments 1 2 5 10 20
Concat 1 48.4 / 64.8
Soup w/ Average 2 49.7 / 66.5 33.9 / 47.4 16.5 / 28.0 11.4 / 22.3
5 46.2 / 60.4 44.0 / 59.1 26.5 / 39.3 17.3 / 29.4
10 16.8 / 27.0 11.1 / 18.3 42.2 / 56.5 34.6 / 47.4
20 19.4 / 30.4 15.2 / 25.4 24.9 / 37.2 24.6 / 38.4